OpenStudy (anonymous):

The 9 officers of the Student Council are going on a trip to an amusement park. Each student must pay an entrance fee plus $5 for meals. The total cost of the trip is $225. Solve the equation 9(e+5)=225 to find the cost e of the entrance fee for each student. A. $20 B. $45 C. $25 D. $14

OpenStudy (ash2326):

We are given the equation \[9(e+5)=225\] 9 represents the 9 officers, e represents the entrance fees for the amusement park and 5 is for the meals. Total cost is 225 we have \[9(e+5)=225\] divide both sides by 9 \[e+5=225/9\] or\[e+5=25\] Subtract 5 form both sides \[e=25-5\] \[e=$20\] so the entrance fees is $20